OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 2) (5, 5) Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 2) (5, 5) @Mathematics

OpenStudy (anonymous):

m in the equation is the slope of your line. To find the slope, you subtract the y components and divide this by the subtracted x components. \[(6-5)/(2-5) = (1/-3) = -1/3\] so your slope is -1/3 b is your y intercept

OpenStudy (anonymous):

first find your slope, which is (y2-y1)/(x2-x1) or (5-2)/(5-6) or -3 this leaves you with the formula y = -3x + b. Plug in one of your known (x,y) and solve for b. 2 = -3(6) + b 2 = -18 + b b = 20 so, now you have your formula. y = -3x + 20